A delivery company's charge for an overnight package weighing in excess of one pound is given by the formula c=2.83w+15.51, where w is the weight of the package in pounds. Find the following.
A 5 pound package cost ___ to send overnight. Please round to the nearest cent.

Answer by Mathtut(3670)

c=2.83w+15.51
:
c(5)=2.83(5)+15.51
:
c(5)=14.15+15.51=$29.66
: